OpenStudy (anonymous):

Please I need help with this question An object moves along a straight line, which we call the x-axis with the positive direction to the right. the equation for the equation for the object's position as a function of time is x(t)= (3+8t-6t^s) m determine A) The average speed from t=3s and t=5s ? I find the answer with negative sign but the speed can not be negative

OpenStudy (anonymous):

1) At t=3 s, x=(3+24+9) m= 36 m (2) At t=5 s, x=(3+40+25) m=68 m as v=(xf−xi)÷(tf−ti)=((68−36)÷(5−3))m/s=16m/s I supposed that the power on t is 2

OpenStudy (vincent-lyon.fr):

x(3s) = -27 m x(5s) = -107 m without any change in direction between the two instants. Finding average velocity is quite easy. Average speed is just the absolute value.